# Forecasting Clinical Risk from Textual Time Series: Structuring Narratives for Temporal AI in Healthcare

### v3

@@ Abstract
Clinical case reports encode temporal patient trajectories that are often underexploited by traditional machine learning methods relying on structured data. In this work, we introduce the forecasting problem from textual time series, where timestamped clinical findings–extracted via an LLM-assisted annotation pipeline–serve as the primary input for prediction. We systematically evaluate a diverse suite of models, including fine-tuned decoder-based large language models and encoder-based transformers, on tasks of event occurrence prediction, temporal ordering, and survival analysis. Our experiments reveal that encoder-based models consistently achieve higher F1 scores and superior temporal concordance for short- and long-horizon event forecasting, while fine-tuned masking approaches enhance ranking performance. In contrast, instruction-tuned decoder models demonstrate a relative advantage in survival analysis, especially in early prognosis settings. Our sensitivity analyses further demonstrate the importance of time ordering, which requires clinical time series construction, as compared to text ordering, the format of the text inputs that LLMs are classically trained on. This highlights the additional benefit that can be ascertained from time-ordered corpora, with implications for temporal tasks in the era of widespread LLM use.
@@ Introduction
Figure 1: Example case report (left) with temporal representations as timelines (middle) and text-ordered, event-time tuples (right).
One might ask whether large pretrained language models can already perform forecasting from clinical narratives, given their exposure to biomedical texts (Peng, Chen, and Lu 2020; Weber et al. 2024). However, these models have architectural limitations for temporal reasoning tasks. Encoder-based models apply random masking that rarely yields temporally coherent representations, while decoder-style models mask text in linear sequence, modeling events in text order rather than time order. Our experiments demonstrate these limitations quantitatively–for example, using the timeline in Figure 1, the probability of recovering a valid time-ordered mask is $0.002\%$, and decoder models achieve low temporal concordance (e.g., $c=0.23$), leading to suboptimal forecasting accuracy without task-specific restructuring.
To address these temporal reasoning challenges, we develop a forecasting framework that transforms free-text narratives into textual time series—sequences of (event, time) tuples and makes explicit the temporal structure of patient trajectories. Building on recent annotation work (Wang and Weiss 2025), we propose a hybrid pipeline combining rule-based heuristics and LLM-assisted extraction to recover temporally localized clinical findings, validated on real clinical notes with expert verification.
@@
In summary, this paper makes the following contributions: (i) Annotation and Extraction Pipeline: Building on prior methods for extracting information from the PubMed Open Access corpus (Noroozizadeh and Weiss 2025), this work introduces a pipeline that transforms case reports into textual time series via regular expression filtering and LLM-assisted extraction, yielding temporally anchored (event, time) tuples while minimizing causal leakage by restricting input to past events; (ii) Comprehensive Model Comparison: We conduct a comprehensive evaluation of encoder- and instruction-tuned decoder-based models—using both fine-tuned MLP heads and prompting—across event prediction, temporal ordering, and survival analysis tasks. Results show no single model excels universally, underscoring the need to align model choice with the specific forecasting objective; (iii) Temporal- versus Text- Ordering: We investigate the impact of annotation order by comparing training on time-ordered versus text-ordered data. Results show that preserving the narrative’s natural order can improve generalization to external datasets, while time-based ordering can enhance ranking performance; (iv) Sensitivity Analysis of Temporal Masking: We conduct systematic dropout experiments by randomly masking parts of the clinical history to assess their impact on forecasting and event ordering. While higher masking levels reduce classification performance (F1), the concordance index remains largely stable, highlighting differing sensitivities of binary prediction and ranking tasks to historical context in textual time series; (v) Methodological Framework for Temporal Clinical Forecasting: Our approach demonstrates how narrative clinical texts can be systematically converted into structured temporal representations for forecasting tasks, providing a replicable methodology that could be adapted to other clinical and non-clinical text sources.
Related Works
@@
Extracting timelines from clinical narratives is a challenging biomedical NLP task. The i2b2 2012 challenge introduced annotations datasets for temporal relation extraction from discharge summaries (Uzuner et al. 2011). Subsequent methods linked clinical events to timestamps or temporal expressions (Leeuwenberg and Moens 2020; Frattallone-Llado et al. 2024), typically using pre-defined event spans. We adopt the approach of Noroozizadeh and Weiss (2025) to assign timepoints to findings from full-length case reports, enabling finer temporal resolution. Unlike methods using structured EHR data, we focus solely on text—crucial for sources like PubMed that lack structured metadata. By directly supervising event-time alignment, we overcome limitations of span-based annotations (Rosenbloom et al. 2011). This aligns with growing emphasis on temporality in sepsis and critical care phenotyping (Johnson et al. 2018; Kamran et al. 2024), particularly given high missingness rates in structured data (Johnson et al. 2023; Seinen et al. 2025).
@@
Clinical text has been used for outcome prediction tasks like mortality and readmission, with models like ClinicalBERT achieving high levels of performance on EHR notes (Huang, Altosaar, and Ranganath 2019; Gu et al. 2021). Text captures complementary information—symptoms and social factors—not found in structured codes. However, traditional approaches using full-text or bag-of-words representations obscure temporal dynamics. Our approach preserves event order within narratives, treating data as time series. This aligns with recent efforts adapting LLMs for time series tasks, such as Time-LLM mapping numeric sequences to tokens (Jin et al. 2023). To our knowledge, this is the first work applying LLMs to text-derived time series in clinical domains. While prior sepsis prediction models rely on vitals or scores like SOFA and SAPS (Hou et al. 2020; Noroozizadeh, Weiss, and Chen 2023), our method is complementary—leveraging narrative descriptions that include clinician interpretations and context. This is valuable when case reports or patient histories are available but structured data is sparse or unavailable.
@@
Methods [^1] Our code is available at https://github.com/Shahriarnz14/Textual-Time-Series-Forecasting
@@
We utilized the PubMed Open Access (PMOA) Subset (as of April 25, 2024) for our analysis. Building on the method described in prior work (Wang and Weiss 2025), the body of each case was extracted by selecting the text between the markers ==== Body and ==== Ref, which serve as section delimiters in the PMOA corpus. To maintain relevance, we include only documents containing case-insensitive matches to (case report|case presenta) and year-? ?old, an approach shown to have higher specificity than PubMed metadata (Noroozizadeh and Weiss 2025).
As demonstrated in prior work of (Noroozizadeh and Weiss 2025), additional filtering was applied to identify potential sepsis-related case reports using the regular expressions sepsi|septic and critical|intensive care. We then used LLM-based queries to extract key attributes, including sepsis diagnosis, patient count, age, and gender. For this step, both GPT-4o and Llama-3.1-8B-Instruct were employed. Reports describing more than one case were excluded. A case report was retained if either model classified it as involving a Sepsis-3 diagnosis.
@@
A textual time series differs from conventional approaches to clinical concept annotation (e.g., (Sun, Rumshisky, and Uzuner 2013; Uzuner et al. 2011)) in its extension of the text span to better capture the specificity and contextual meaning of each clinical finding. In our work, the interpretation of a clinical finding differs from the i2b2 concept guidelines in the following respects : (1) Clinical findings are not restricted to single prepositional phrases following a markable entity. For instance, instead of splitting “pain in chest that radiates substernally” into less informative parts, we retain the full phrase as a single, complete finding to preserve its meaning and context. (2) To improve clarity, coordinated phrases are broken into individual findings. For example, “metastases in the liver and pancreas” becomes “metastasis in the liver” and “metastasis in the pancreas.”
We used DeepSeek-R1-UD-IQ1 and Llama-3.3-70B-Instruct models to generate clinical textual time series from the 2319 sepsis case reports. We refer to this as Textual Time-Series for Sepsis (T2S2). The exact LLM extraction prompt can be found in Appendix E. Our dataset was split into training (n = 1,795; 80% train, 20% validation), testing T2S2-test (n = 244), and two external validation sets: sepsis-10 (n = 10) and sepsis-100 (n = 90).
@@
This task assesses the model’s ability to reconstruct the correct sequence of future events. At each time cut-off $t$, we extract the next $k$ events from the timeline and remove their timestamps. The model must correctly output a permutation over these events that matches their true chronological order. This is framed as a ranking task, evaluated by computing the pairwise concordance between the predicted and true orderings (e.g., proportion of correctly ordered pairs). This tests whether the model can infer temporal progression from unordered event content alone.
@@
We include a classical survival analysis task to model time until death. Many case reports specify whether the patient died and, if so, when (e.g., ”the patient died on hospital day 10”), which enables us to define a time-to-event outcome. For this task, we evaluate models at predefined cut-off times – specifically at 0 hours (admission), 24 hours (1 day), and 168 hours (1 week) – and use the extracted event history up to each cut-off as input. A survival model is trained to predict the probability of survival over time beyond each cut-off. We evaluate using the time-dependent concordance index to measure alignment of the predicted survival times with actual outcomes.
@@
To evaluate performance on the event forecasting and survival prediction tasks, we implement five modeling paradigms: (i) decoder-only large language models (LLMs) with fine-tuned heads, (ii) prompted LLMs without gradient updates (zero- or few-shot), (iii) encoder-only models with task-specific fine-tuned heads, (iv) encoder-masking models with fine-tuning, and (v) encoder-masking models in zero-shot settings. In Appendix C and  D, we provide detailed explanations of implementation details.
@@
We apply instruction-tuned decoder-only models from the LLaMA and DeepSeek families: Llama-3.3-70B-Instruct, Llama-3.1-8B-Instruct, and their distilled variants (Grattafiori et al. 2024; Liu et al. 2024). We also include open-source models with documented training corpora (OLMO-32B-Instruct (OLMo et al. 2024), RedPajama-INCITE-7B-Instruct (Weber et al. 2024)) to address potential data leakage concerns–notably, RedPajama explicitly excludes PubMed from its training data. Additionally, we evaluate a medically fine-tuned decoder model (MediPhi-PubMed (Corbeil et al. 2025)) to assess domain-specific benefits.
@@
In the zero- or few-shot setting, we use the same LLM architectures as above, but without any fine-tuning. Instead, we supply structured prompts at inference time to guide the model toward task-specific outputs. Each prompt includes: (1) a system instruction establishing the model’s role (e.g., ”You are an expert physician.”), (2) a user instruction describing the prediction task, and (3) one or more input-output examples in a few-shot format. Prompt designs are customized for each task and are detailed in Appendix E. Model outputs are parsed to extract binary labels or ordered lists, and errors in generation or ambiguity are excluded from evaluation.
@@
We conduct sensitivity analyses to examine how the temporal structure and completeness of input sequences affect forecasting performance. These analyses probe the robustness of models to changes in event ordering and to varying levels of missing historical information:
@@
We evaluate two strategies for ordering clinical events within textual time series inputs. In the text-ordered setting, events are presented in the order they appear in the original case report narratives, preserving the narrative structure produced during extraction. In the time-ordered setting, events are sorted chronologically by their documented occurrence time, enforcing strict temporal alignment. This comparison isolates the effect of narrative sequencing versus explicit temporal ordering on model performance. For text-ordered inputs, to prevent causal leakage, events occurring after the forecast cutoff time $t$ are masked for encoder-based models and omitted for decoder-based models.
@@
Together, these analyses characterize model robustness to disruptions in temporal coherence and input completeness – two key challenges in real-world deployment scenarios.
Table 1: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events. All models are trained/fine-tuned on time-ordered annotations from either DeepSeek-R1 or Llama 3.3-70B. Bold values indicate best in category within each column group (refer to Tables A1 and A2 in Appendix A for detailed performance statistics on all models). Performance statistics for all models are presented in A1 and A2, respectively. Abbreviations: F1(1d)-10/100: F1 (1 day) for sepsis-10 and sepsis-100 respectively. c10/100: c-index for sepsis-10 and sepsis-100 respectively.
T2S2 (DeepSeek-R1) T2S2 (Llama 3.3-70B)
@@ Results
In this section, we present results for our three main evaluation settings: event forecast within a subsequent time window, temporal ordering of forecasted events, and survival analysis. We then provide sensitivity analyses on the forecasting tasks, examining the effects of temporal ordering and historical context availability.
@@
Our results show that encoder-based models outperform decoder-based LLMs in event forecasting across both DeepSeek-R1 and Llama-3.3-70B annotations (Table 1). Forecasting performance across all models is shown in Tables A1 and A2 (Appendix A). Encoder models, especially those with a fine-tuned MLP head, achieve substantially higher F1 scores than decoder LLMs, reinforcing the effectiveness of encoder-based representations for forecasting.
Among decoder models, we observe varied performance patterns. RedPajama-INCITE-7B-Instruct, with documented training that excludes PubMed, achieves 0.352 F1 at 24h—better than some decoder models like MediPhi-PubMed (0.268) and within the range of decoder performance, though below top performers like DeepSeek-Llama-3.3-70B (0.482). OLMO-32B-Instruct performs competitively (0.411), suggesting that open models without biomedical pretraining can still achieve reasonable results for temporal forecasting tasks.
@@
Encoder models consistently outperform decoders in correctly ranking the order of upcoming events, as measured by concordance (c-index; Table 1; complete results in Tables A1 and A2 for DeepSeek-R1 and Llama-3.3-70B annotations, respectively, with full details in Appendix A. Decoder models with potential biomedical exposure achieve modest concordance (0.618-0.632), while RedPajama-INCITE-7B-Instruct without PubMed data in its training corpus performs similarly, suggesting domain-specific knowledge is less critical here than for event prediction. The comparable performance of RedPajama-INCITE-7B-Instruct to other general models (with access to PubMed in their pretraining) also provides evidence that the T2S2 prediction task is distinct from raw PubMed text and assuages causal leakage concerns.
Biomedical pretraining in encoder models also demonstrates clear advantages, with BioClinical-ModernBERT-base attaining the highest concordance (0.677 with MLP head, 0.676 fine-tuned masking) and robust performance across datasets. MLP-head training yields superior F1 scores, while fine-tuned masking excels in concordance. BioClinical-ModernBERT-base maintains strong generalization on external datasets (c-index >0.60), suggesting that biomedical pretraining provides stability in temporal reasoning across dataset shifts.
@@
Table 2 presents time-dependent concordance across evaluation cohorts (T2S2, sepsis-10, sepsis-100) at multiple observation windows (0h, 24h, 168h), selecting models based on validation concordance. General models, such as RedPajama-INCITE-7B-Instruct, show strong survival prediction capabilities even without biomedical exposure (0.76 concordance at 168h, sepsis-100). Meanwhile, biomedical encoders like BioClinical-ModernBERT-base consistently perform well, particularly early (0h).
Predictive performance does not universally improve with more observation time. T2S2 cohort models often peak at earlier windows, while sepsis-10 experiences ceiling effects. Instruction-tuned decoders excel overall, yet biomedical encoders remain competitive, particularly at admission (0h). Survival prediction thus benefits from both larger decoder models and domain-specific pretraining, each excelling in different temporal contexts.
@@
Table 2: Survival analyses evaluated using time-dependent concordance index for each model across DeepSeek-R1 test set annotations (T2S2, Sepsis-10, Sepsis-100), evaluated at three observation time thresholds (0h, 24h, and 168h).
Model T2S2 sepsis-10 sepsis-100
Time Threshold 0h 24h 168h 0h 24h 168h 0h 24h 168h
bert-base-uncased 0.60 0.61 0.53 1.00 1.00 1.00 0.58 0.62 0.58
roberta-base 0.55 0.60 0.60 0.75 1.00 1.00 0.55 0.58 0.57
deberta-v3-small 0.56 0.57 0.57 0.75 0.00 0.75 0.52 0.49 0.51
ModernBERT-base 0.52 0.58 0.58 1.00 1.00 1.00 0.54 0.51 0.51
ModernBERT-large 0.53 0.59 0.58 0.50 1.00 1.00 0.59 0.55 0.49
DeepSeek-R1-Llama-70B 0.64 0.63 0.59 0.50 1.00 1.00 0.61 0.53 0.55
Llama-3.3-70B-Instruct 0.62 0.63 0.58 0.50 1.00 0.75 0.64 0.57 0.53
DeepSeek-R1-Llama-8B 0.60 0.58 0.58 1.00 1.00 1.00 0.53 0.66 0.65
Llama-3.1-8B-Instruct 0.62 0.61 0.61 1.00 1.00 1.00 0.60 0.59 0.61
OLMO 32B Instruct 0.60 0.62 0.59 0.75 0.75 0.75 0.66 0.58 0.56
MediPhi-PubMed 0.54 0.62 0.55 1.00 1.00 1.00 0.62 0.58 0.61
Sensitivity analysis
Time-Ordered versus Text-Ordered Training.
Table 3 shows time-ordering generally improves concordance, particularly for DeepSeek-R1 annotations, while F1 scores show mixed results. Text-ordering sometimes yields better external dataset performance with Llama-3.3-70B, suggesting that preserving the narrative sequence can support accurate classification even when it does not follow temporal order. These results highlight a trade-off between temporal alignment and narrative coherence for different tasks.
Table 3: Comparison of text-ordered and time-ordered training for DeepSeek-R1 and Llama 3.3-70B annotations on F1 scores and c-index values.
F1(1h) F1(1d) F1(1w) c-index F1(1d)-10 c10 F1(1d)-100 c100
T2S2 (DeepSeek-R1)
Text-ordered 0.290 0.569 0.876 0.659 0.361 0.598 0.615 0.609
Time-ordered 0.306 0.576 0.877 0.672 0.428 0.576 0.607 0.613
T2S2 (Llama 3.3-70B)
Text-ordered 0.311 0.648 0.898 0.586 0.423 0.526 0.662 0.541
Time-ordered 0.257 0.645 0.903 0.632 0.382 0.595 0.616 0.612
Impact of timestep dropout on F1 and concordance.
Tables 4 and A3 (Appendix B) report the effect of timestep dropout on classification and ranking performance. F1 scores decrease monotonically with increasing TDR, with relatively mild degradation up to 30–60% dropout and sharper declines thereafter. For example, in the DeepSeek-R1 annotations, T2S2/sepsis-10/sepsis-100 drop from 0.576/0.428/0.607 at TDR = 0% to 0.423/0.200/0.510 at TDR = 90%, with sepsis-10 showing the largest relative decline. In contrast, concordance (c-index) remains stable across dropout levels (0.661–0.678), indicating that ranking-based evaluations are less sensitive to history truncation. These trends are consistent for Llama-3.3 annotations (Appendix B), suggesting that ranking objectives capture temporal relationships that are more robust to missing historical information than binary classification accuracy of predicting whether an events happens within a time window.
Table 4: Effect of Randomly Masking History (DeepSeek-R1 annotations) on F1 and Concordance.
Timestep drop rate F1(1d) c-index F1(1d)-10 c10 F1(1d)-100 c100
0% 0.576 0.672 0.428 0.576 0.607 0.613
15% 0.572 0.667 0.415 0.564 0.604 0.616
30% 0.548 0.672 0.357 0.542 0.591 0.619
60% 0.542 0.678 0.310 0.578 0.588 0.616
90% 0.423 0.661 0.200 0.561 0.510 0.614
Discussion

### v5

@@ Abstract
Clinical case reports encode temporal patient trajectories that are often underexploited by traditional machine learning methods relying on structured data. In this work, we introduce the forecasting problem from textual time series, where timestamped clinical findings—extracted via an LLM-assisted annotation pipeline—serve as the primary input for prediction. We systematically evaluate a diverse suite of models, including fine-tuned decoder-based large language models and encoder-based transformers, on tasks of event occurrence prediction, temporal ordering, and survival analysis. Our experiments reveal that encoder-based models consistently achieve higher F1 scores and superior temporal concordance for short- and long-horizon event forecasting, while fine-tuned masking approaches enhance ranking performance. In contrast, instruction-tuned decoder models demonstrate a relative advantage in survival analysis, especially in early prognosis settings. Our sensitivity analyses further demonstrate the importance of time ordering, which requires clinical time series construction, as compared to text ordering, the format of the text inputs that LLMs are classically trained on. This highlights the additional benefit that can be ascertained from time-ordered corpora, with implications for temporal tasks in the era of widespread LLM use.
@@ Introduction
Figure 1: Example case report (left) with timeline representations (middle) and text-ordered, (event, time) tuples (right).
One might ask whether large pretrained language models can already perform forecasting from clinical narratives, given their exposure to biomedical texts (Peng, Chen, and Lu 2020; Weber et al. 2024). However, these models have architectural limitations for temporal reasoning tasks. Encoder-based models apply random masking that rarely yields temporally coherent representations, while decoder-style models mask text in linear sequence, modeling events in text order rather than time order. Our experiments make these limitations explicit. For instance, on the timeline in Figure 1, the chance of recovering a valid time-ordered mask is below $0.002\%$, and decoder models show low temporal concordance (e.g., $c=0.23$). Without a task-specific restructuring, such factors lead to suboptimal forecasting performance.
To address these temporal reasoning challenges, we develop a forecasting framework that transforms free-text narratives into textual time series—sequences of (event, time) tuples—and makes explicit the temporal structure of patient trajectories. Building on recent annotation work (Wang and Weiss 2025; Noroozizadeh et al. 2025), we propose a hybrid pipeline combining rule-based heuristics and LLM-assisted extraction to recover temporally localized clinical findings, validated on real clinical notes with expert verification.
@@
In summary, this paper makes the following contributions:
(i) Annotation and Extraction Pipeline: Building on prior methods for extracting information from the PubMed Open Access corpus (Noroozizadeh and Weiss 2025), this work introduces a pipeline that transforms case reports into textual time series via regular expression filtering and LLM-assisted extraction, yielding temporally anchored (event, time) tuples while minimizing causal leakage by restricting input to past events; (ii) Comprehensive Model Comparison: We conduct a comprehensive evaluation of encoder- and instruction-tuned decoder-based models—using both fine-tuned MLP heads and prompting—across event prediction, temporal ordering, and survival analysis tasks. Results show no single model excels universally, underscoring the need to align model choice with the specific forecasting objective; (iii) Temporal- versus Text- Ordering: We investigate the impact of annotation order by comparing training on time-ordered versus text-ordered data. Results show that preserving the narrative’s natural order can improve generalization to external datasets, while time-based ordering can enhance ranking performance; (iv) Sensitivity Analysis of Temporal Masking: We conduct systematic dropout experiments by randomly masking parts of the clinical history to assess their impact on forecasting and event ordering. While higher masking levels reduce classification performance (F1), the concordance index remains largely stable, highlighting differing sensitivities of binary prediction and ranking tasks to historical context in textual time series; (v) Methodological Framework for Temporal Clinical Forecasting: Our approach demonstrates how narrative clinical texts can be systematically converted into structured temporal representations for forecasting tasks, providing a replicable methodology that could be adapted to other clinical and non-clinical text sources.
Related Work
@@
Extracting timelines from clinical narratives is a challenging biomedical NLP task. The i2b2 2012 challenge introduced annotated datasets for temporal relation extraction from discharge summaries (Uzuner et al. 2011). Subsequent methods linked clinical events to timestamps or temporal expressions (Leeuwenberg and Moens 2020; Frattallone-Llado et al. 2024), typically assuming pre-defined event spans. We adopt our prior approach (Noroozizadeh and Weiss 2025; Noroozizadeh et al. 2025) to assign timepoints to findings from full-length case reports, enabling finer temporal resolution. Unlike methods using structured EHR data, we focus solely on text—crucial for sources like PubMed that lack structured metadata. By directly supervising event-time alignment, we overcome limitations of span-based annotations (Rosenbloom et al. 2011). This aligns with growing emphasis on temporality in sepsis and critical care phenotyping (Johnson et al. 2018; Kamran et al. 2024), particularly given high missingness rates in structured data (Johnson et al. 2023; Seinen et al. 2025).
@@
Clinical text has been used for outcome prediction tasks like mortality and readmission, with models like ClinicalBERT achieving high levels of performance on EHR notes (Huang, Altosaar, and Ranganath 2019; Gu et al. 2021). Text captures complementary information—symptoms and social factors—not found in structured codes. However, traditional approaches using full-text or bag-of-words representations obscure temporal dynamics. Our approach preserves event order within narratives, treating data as time series. This aligns with recent efforts adapting LLMs for time series tasks, such as Time-LLM mapping numeric sequences to tokens (Jin et al. 2023). To our knowledge, this is the first work to forecast from LLM-derived time series in clinical domains. While prior sepsis prediction models rely on vitals or scores like SOFA and SAPS (Hou et al. 2020; Noroozizadeh, Weiss, and Chen 2023), our method is complementary—leveraging narrative descriptions that include clinician interpretations and context. This is valuable when case reports or patient histories are available but structured data is sparse or unavailable.
@@
Methods [^1] Our code can be found at: https://github.com/Shahriarnz14/Textual-Time-Series-Forecasting.
   The extended version of this paper, including the full technical appendix, is available in Noroozizadeh, Kumar, and Weiss (2025).
@@
We used the PubMed Open Access (PMOA) Subset (as of April 25, 2024) for our analysis. Following prior work (Wang and Weiss 2025), we extracted the body of each case by selecting the text between the ==== Body and ==== Ref delimiters in the PMOA corpus. To maintain relevance, we included only documents with case-insensitive matches to (case report|case presenta) and year-? ?old, an approach shown to be more specific than relying on PubMed metadata (Noroozizadeh and Weiss 2025).
Additional filtering was applied to identify potential sepsis-related case reports using the regular expressions sepsi|septic and critical|intensive care. We then used LLM-based queries to extract key attributes, including sepsis diagnosis, patient count, age, and gender. For this step, both GPT-4o and Llama-3.1-8B-Instruct were employed. Reports describing more than one case were excluded. A case report was retained if either model classified it as involving a Sepsis-3 diagnosis.
@@
A textual time series differs from conventional approaches to clinical concept annotation (e.g., (Sun, Rumshisky, and Uzuner 2013; Uzuner et al. 2011)) in its extension of the text span to better capture the specificity and contextual meaning of each clinical finding. In our work, the interpretation of a clinical finding differs from the i2b2 concept guidelines in the following respects: (1) Clinical findings are not restricted to single prepositional phrases following a markable entity. For instance, instead of splitting “pain in chest that radiates substernally” into less informative parts, we retain the full phrase as a single, complete finding to preserve its meaning and context. (2) To improve clarity, compound phrases are broken into individual findings. For example, “metastases in the liver and pancreas” becomes “metastasis in the liver” and “metastasis in the pancreas.”
We used DeepSeek-R1-UD-IQ1 and Llama-3.3-70B-Instruct models to generate clinical textual time series from the 2319 sepsis case reports. We refer to this as Textual Time-Series for Sepsis (T2S2). Details on the quality evaluation of LLM-extracted temporal annotations and the prompts used for extraction are provided in Appendix A and A, respectively. Our dataset was split into training (n = 1,795; 80% train, 20% validation), testing T2S2-test (n = 244), and two external validation sets: sepsis-10 (n = 10) and sepsis-100 (n = 90).
@@
This task assesses the model’s ability to reconstruct the correct sequence of future events. At each time cut-off $t$, we extract the next $k$ events from the timeline and remove their timestamps. The model must output a permutation of these events that matches their true chronological order. This is framed as a ranking task, evaluated by computing the pairwise concordance between the predicted and true orderings (e.g., proportion of correctly ordered pairs). This tests whether the model can infer temporal progression from unordered event content alone.
@@
We include a classical survival analysis task to model time until death. Many case reports specify whether the patient died and, if so, when (e.g., ”the patient died on hospital day 10”), which enables us to define a time-to-event outcome. For this task, we evaluate models at predefined cut-off times— specifically at 0 hours (admission), 24 hours (1 day), and 168 hours (1 week)— and use the extracted event history up to each cut-off as input. A survival model is trained to predict the probability of survival over time beyond each cut-off. We evaluate using the time-dependent concordance to measure alignment of the predicted survival times with actual outcomes.
@@
To evaluate performance on the event forecasting and survival prediction tasks, we implement five modeling paradigms: (i) decoder-only large language models (LLMs) with fine-tuned heads, (ii) prompted LLMs without gradient updates (zero- or few-shot), (iii) encoder-only models with task-specific fine-tuned heads, (iv) encoder-masking models with fine-tuning, and (v) encoder-masking models in zero-shot settings. In Appendix D and E, we provide technical details of encoder-based forecasting methods and the survival modeling framework respectively.
@@
We apply instruction-tuned decoder-only models from the Llama and DeepSeek families: Llama-3.3-70B-Instruct, Llama-3.1-8B-Instruct, and their distilled variants (Grattafiori et al. 2024; Liu et al. 2024a). We also include open-source models with documented training corpora (OLMO-32B-Instruct (OLMo et al. 2024), RedPajama-INCITE-7B-Instruct (Weber et al. 2024)) to address potential data leakage concerns—notably, RedPajama explicitly excludes PubMed from its training data. Additionally, we evaluate a medically fine-tuned decoder model (MediPhi-PubMed (Corbeil et al. 2025)) to assess domain-specific benefits.
@@
In the zero- or few-shot setting, we use the same LLM architectures as above, but without any fine-tuning. Instead, we supply structured prompts at inference time to guide the model toward task-specific outputs. Each prompt includes: (1) a system instruction establishing the model’s role (e.g., ”You are an expert physician.”), (2) a user instruction describing the prediction task, and (3) one or more input-output examples in a few-shot format. Prompt designs are customized for each task and are detailed in Appendix A. Model outputs are parsed to extract binary labels or ordered lists, and errors in generation or ambiguity are excluded from evaluation.
@@
We conduct sensitivity analyses to examine how the temporal structure and completeness of input sequences affect forecasting performance. These analyses probe the robustness of models to changes in event ordering and to varying levels of missing historical information.
@@
We evaluate two strategies for ordering clinical events within textual time series inputs. In the text-ordered setting, events are presented in the order they appear in the original case report narratives, preserving the narrative structure produced during extraction. In the time-ordered setting, events are sorted chronologically by their documented occurrence time, enforcing strict temporal alignment. This comparison isolates the effect of narrative sequencing versus explicit temporal ordering on model performance. For text-ordered inputs, to prevent causal leakage, events occurring after the forecast cutoff time $t$ are masked for encoder-based models and omitted for decoder models.
@@
T2S2 (DeepSeek-R1) T2S2 (Llama-3.3-70B)
@@
Table 1: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events. All models are trained/fine-tuned on time-ordered annotations from either DeepSeek-R1 or Llama 3.3-70B. Bold values indicate best in category within each column group (refer to Tables A3 and A4 in Appendix E for detailed performance statistics on all models). Performance statistics for all models are presented in A3 and A4, respectively. Abbreviations: F1(1d)-10/100: F1 (1 day) for sepsis-10 and sepsis-100 respectively. c10/100: c-index for sepsis-10 and sepsis-100 respectively.
@@ Results
In this section, we present results for our three main evaluation settings: event forecast within a subsequent time window, temporal ordering of forecasted events, and survival analysis. We provide sensitivity analyses on the forecasting tasks, examining the effects of temporal ordering and historical context availability in Appendix E and E.
@@
Our results show that encoder-based models outperform decoder-based LLMs in event forecasting across both DeepSeek-R1 and Llama-3.3-70B annotations (Table 1). Forecasting performance across all models is shown in Tables A3 and A4 (Appendix E). Encoder models, especially those with a fine-tuned MLP head, achieve substantially higher F1 scores than decoder LLMs, reinforcing the effectiveness of encoder-based representations for forecasting.
Among decoders, we observe varied performance patterns. RedPajama-INCITE-7B-Instruct, whose training excludes PubMed, achieves 0.352 F1 at 24h—better than some decoder models like MediPhi-PubMed (0.268) and within the typical decoder performance range, though below top performers like DeepSeek-Llama-3.3-70B (0.482). OLMO-32B-Instruct performs competitively (0.411), suggesting that open models without biomedical pretraining can still achieve strong temporal forecasting results.
@@
Encoder models consistently outperform decoders in correctly ranking the order of upcoming events, as measured by concordance (c-index; Table 1; complete results in Tables A3 and A4 for DeepSeek-R1 and Llama-3.3-70B annotations, respectively, with full details in Appendix E. Decoder models with potential biomedical exposure achieve modest concordance (0.618-0.632), while RedPajama-INCITE-7B-Instruct without PubMed data in its training corpus performs similarly, suggesting domain-specific knowledge is less critical here than for event prediction. The comparable performance of RedPajama-INCITE-7B-Instruct to other general models (with access to PubMed in their pretraining) also provides evidence that the T2S2 prediction task is distinct from raw PubMed text and assuages causal leakage concerns.
Biomedical pretraining in encoders demonstrates clear advantages, with BioClinical-ModernBERT-base achieving the highest concordance (0.677 with an MLP head, 0.676 with fine-tuned masking) and strong performance across datasets. MLP-head training yields superior F1 scores, while fine-tuned masking excels in concordance. BioClinical-ModernBERT-base also generalizes well to external datasets (c-index >0.60), suggesting that biomedical pretraining provides stability in temporal reasoning under dataset shifts.
@@
Tables 2 and A5 (Appendix E) show time-dependent concordance results where instruction-tuned decoder models clearly outperform encoder-based models. This pattern contrasts with the forecasting tasks (Table 1), where encoders excelled. Across both the T2S2 and sepsis-100 cohorts, decoder models consistently achieve the strongest results. Notably, RedPajama-INCITE-7B-Instruct attains the highest concordance overall (0.76 at 168h, sepsis-100) despite lacking any biomedical text in pretraining, while larger models such as DeepSeek-R1-Llama-70B and OLMO-32B-Instruct also perform robustly. Encoder models—including biomedically fine-tuned variants like BioClinical-ModernBERT-base, which is competitive on sepsis-100 but not on T2S2—are generally surpassed by decoder models.
Predictive performance does not universally improve with more observation time. Models evaluated on the T2S2 cohort, often peak at earlier windows (0h or 24h) rather than at 168h. Meanwhile, the sepsis-10 cohort experiences clear ceiling effects, with many models achieving perfect concordance (1.00). These results suggest that, unlike the forecasting tasks, survival prediction in our dataset benefits more from the capabilities inherent in large decoder models than from domain-specific pretrained encoders.
Model 0h 24h 168h
bert-base-uncased 0.60 0.61 0.53
roberta-base 0.55 0.60 0.60
deberta-v3-small 0.56 0.57 0.57
ModernBERT-base 0.52 0.58 0.58
ModernBERT-large 0.53 0.59 0.58
Bioclinical ModernBERT-base 0.57 0.53 0.53
Bioclinical ModernBERT-large 0.54 0.55 0.54
DeepSeek-R1-Llama-70B 0.64 0.63 0.59
Llama-3.3-70B-Instruct 0.62 0.63 0.58
DeepSeek-R1-Llama-8B 0.60 0.58 0.58
Llama-3.1-8B-Instruct 0.62 0.61 0.61
OLMO 32B Instruct 0.60 0.62 0.59
RedPajama-INCITE 7B Instruct 0.56 0.60 0.61
MediPhi-PubMed 0.54 0.62 0.55
@@
Table 2: Time-dependent concordance index for survival analysis on the DeepSeek-R1 T2S2 annotations evaluated at 0h, 24h, and 168h. Extended sepsis-10 and sepsis-100 results are provided in Appendix E.
Discussion and Conclusion